OpenStudy (sreekar369):

A pen costs 13 rupees and a note book costs 35 rupees.let m be the maximum number of items that can be bought for 1000 rupees and n be the minimum number of items that can be bought for the same amount ,then m+n is a)76 b)88 c)96 d)98

OpenStudy (sreekar369):

@sumanta

OpenStudy (sreekar369):

will give a medal

OpenStudy (sreekar369):

@Michele_Laino

OpenStudy (michele_laino):

@sreekar369 I have found!, it is m+n=76

OpenStudy (michele_laino):

Please you have to solve this equation: \[13*x+35*y=1000\] where x is the number of pens, and y is the number of notebook. Now, I rewrite the above equation in the new form: \[35*y=1000-13*x\] that equation, tells us that the difference between 1000 and 35*x must be divisible for 35, so I try to set x=5: substituting, we have: \[35*y=1000-55=945\] from which I get y945/35=27, finally I can write x+y=27+5=32=n. Next step is to set x=20, so we have to solve the subsequent equation: \[35*y=1000-260=840\] from which I get y= 840/35=24, so x+y=20+24=44=m. FInally I have m+n=32+44=76.

OpenStudy (sreekar369):

thqs

OpenStudy (michele_laino):

thanks!